write the equation of the line that passes through (0,-4) and is perpendicular to x-3y=15, in standard form
Answer by mananth(16946)

1 x + -3 y = 15
Find the slope of this line

-3 y = -1 x + 15
Divide by -3
y = 1/ 3 x + -5
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
Because m1*m2 =-1
The slope of the required line will be -3

m= -3 ,point ( 0 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = -0 + b
b= -4
m= -3
The required equation is y = -3 x -4
y+3x=-4
m.ananth@hotmail.ca